Steve races to the nearest taco stand at lunchtime and sees that his pedometer recorded his peak speed at 76.7 cm/s. What was Steve's peak speed in kilometers per hour?
1
Expert's answer
2019-09-25T05:43:12-0400

"1\ Km=100\times 1000\ cm\\1\ hour=60\times 60=3600\ sec"

Now

"1" "\ cm=10^{-5}\ km\\"

"1\ sec=0.00027777777
\ hour"

Now

"speed=76.7\frac{\ cm}{sec}=\frac{76.7\times{10^{-5}}}{0.00027777777}\frac{\ km}{hour}=2.7612\ km\ hour^{-1}"